OpenStudy (anonymous):

find two consecutive numbers whose product is 29 more than their sum.

OpenStudy (anonymous):

Note that if n is an integer, then n and n+1 are consecutive integers. What you want to do is find n so that \(\large n(n+1) = 29 + n+(n+1)\). Can you take things from here? :-)

OpenStudy (anonymous):

(X)(X+1)=29+(X)+(X+1) X^2 + X = 30 + 2X X^2 -X -30 = 0 (X-6)(X+5)=0 X=6, X=-5